# Integrated Sensing and Communications for Pinching-Antenna Systems (PASS)

### v1

@@ III-B Antenna Activation Position Optimization
To deal with the intractable expression of variable channels, we introduce auxiliary variables $\mathbf{U}_{1}$, $\mathbf{V}_{1}$, $\mathbf{U}_{2}$, and $\mathbf{V}_{2}$ to replace $\mathbf{H}_{1,\text{s}}$, $\mathbf{H}_{1,\text{c}}$, $\mathbf{H}_{2,\text{s}}$, and $\mathbf{H}_{2,\text{c}}$, respectively. Thus, we have the equality constraints $\mathbf{U}_{1}=\mathbf{H}_{1,\text{s}}$, $\mathbf{V}_{1}=\mathbf{H}_{2,\text{s}}$, $\mathbf{U}_{2}=\mathbf{H}_{1,\text{c}}$, and $\mathbf{V}_{2}=\mathbf{H}_{2,\text{c}}$. By relocating the equality constraint to the objective function and serving as a penalty term, the problem (P3) can be equivalently rewritten as
$(\text{P}5)\quad$ $\max\limits_{\mathbf{x}^{\text{p}},\mathbf{U}_{i},\mathbf{V}_{i}}\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (12a)
s.t. $\text{Tr}(\mathbf{U}_{2}\mathbf{W})\geq\gamma_{\text{QoS}}[\text{Tr}(\mathbf{V}_{2}\mathbf{R}_{\text{s}})+\sigma^{2}],$ (12b)
@@
$\frac{1}{r_{\text{min,s}}^{2}}\leq\frac{\mathbf{X}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,s}}^{2}},\ \mathbf{X}\in\{\mathbf{U}_{1},\mathbf{V}_{1}\},$ (12e)
$\frac{1}{r_{\text{min,c}}^{2}}\leq\frac{\mathbf{Y}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,c}}^{2}},\ \mathbf{Y}\in\{\mathbf{U}_{2},\mathbf{V}_{2}\},$ (12f)
@@
where $\chi=\|\mathbf{U}_{1}-\mathbf{H}_{1,\text{s}}\|_{F}+\|\mathbf{V}_{1}-\mathbf{H}_{2,\text{s}}\|_{F}+\|\mathbf{U}_{2}-\mathbf{H}_{1,\text{c}}\|_{F}+\|\mathbf{V}_{2}-\mathbf{H}_{2,\text{c}}\|_{F}$, and $\varrho$ denotes the scaling factor of the penalty terms. Here, $r_{\text{min,s}}$ and $r_{\text{max,s}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the sensing target. Also, $r_{\text{min,c}}$ and $r_{\text{max,c}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the communication user. To solve the problem (P5), we propose a penalty-based two-layer algorithm, which alternately optimizes $\{\mathbf{U},\mathbf{V}\}$ and $\{\mathbf{x}^{\text{p}}\}$ in the inner layer and update $\varrho$ in the outer layer.
@@ III-B1 Inner layer iteration—subproblem with respect to $\{\mathbf{U},\mathbf{V}\}$
$(\text{P}6)\ \max\limits_{\mathbf{x}^{\text{p}},\mathbf{U}_{i},\mathbf{V}_{i}\atop\mathbf{W},\mathbf{R}_{\text{s}}}$ $\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (13a)
@@
$\begin{cases}\Re(\text{Tr}(\mathbf{U}_{i}^{H}(\mathbf{I}-\mathbf{u}_{i}\mathbf{u}_{i}^{H})))\leq\varrho_{1},\\ \Re(\text{Tr}(\mathbf{V}_{i}^{H}(\mathbf{I}-\mathbf{v}_{i}\mathbf{v}_{i}^{H})))\leq\varrho_{2},\end{cases}\ i\in\{1,2\},$ (14)
where $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$ represent the leading eigenvectors of $\mathbf{U}_{i}$ and $\mathbf{V}_{i}$. As a result, the problem (P6) can be transformed into
$(\text{P}7)\ \max\limits_{\mathbf{U}_{i},\mathbf{V}_{i},\varrho_{i}}\quad$ $\text{Tr}(\mathbf{U}_{1}\mathbf{W})+\text{Tr}(\mathbf{V}_{1}\mathbf{R}_{\text{s}})-\frac{1}{2\varrho}\chi$ (15a)
@@
which is a convex problem and can be directly solved. Thus, the rank-one solution $\{\mathbf{U}_{i},\mathbf{V}_{i}\}$ can be obtained by iteratively solving the problem (P7).
@@ III-B2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
Note that the equality constraint $\mathbf{U}_{1}=\mathbf{H}_{1,\text{s}}$, $\mathbf{V}_{1}=\mathbf{H}_{2,\text{s}}$, $\mathbf{U}_{2}=\mathbf{H}_{1,\text{c}}$ are equivalent to $\mathbf{u}_{1}=\mathbf{h}_{1,\text{s}}$, $\mathbf{v}_{1}=\mathbf{h}_{2,\text{s}}$, $\mathbf{u}_{2}=\mathbf{h}_{1,\text{c}}$ and $\mathbf{v}_{2}=\mathbf{h}_{2,\text{c}}$, where $\{\mathbf{u}_{1},\mathbf{v}_{1},\mathbf{u}_{2},\mathbf{v}_{2}\}$ denote the leading eigenvectors of $\{\mathbf{U}_{1},\mathbf{V}_{1},\mathbf{U}_{2},\mathbf{V}_{2}\}$. As a result, the problem (P6) can be transformed into
$(\text{P}8)\ \min\limits_{\mathbf{x}^{\text{p}}}\quad$ $\|\mathbf{u}_{1}-\mathbf{h}_{1,\text{s}}\|^{2}+\|\mathbf{v}_{1}-\mathbf{h}_{2,\text{s}}\|^{2}+\|\mathbf{u}_{2}-\mathbf{h}_{1,\text{c}}\|^{2}$
$+\|\mathbf{v}_{2}-\mathbf{h}_{2,\text{c}}\|^{2}$ (16a)
@@
It is easy to notice that $x_{n}^{\text{p}}$ and $x_{m}^{\text{p}}$ ($n\neq m$) are separated in the objective function but coupled in the constraint (8h), which motivates us to adopt the element-wise optimization framework. Therefore, with the fixed $\{x_{1}^{\text{p}},\cdots,x_{n-1}^{\text{p}},x_{n+1}^{\text{p}},\cdots,x_{N}^{\text{p}}\}$, the subproblem with respect to $x_{n}^{\text{p}}$ is given by
@@
$\varpi(x_{n}^{\text{p}})=\begin{cases}\|\mathbf{u}_{[n],1}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{u}_{[n],2}-\mathbf{h}_{[n],\text{c}}\|^{2},\ n\in\mathcal{N}_{1},\\ \|\mathbf{v}_{[n],1}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{v}_{[n],2}-\mathbf{h}_{[n],\text{c}}\|^{2},\ n\in\mathcal{N}_{2}.\end{cases}$ (18)
where $\mathcal{N}_{1}=\{1,\cdots,N_{1}\}$ and $\mathcal{N}_{2}=\{N_{1}+1,\cdots,N\}$. Then, the optimal $x_{n}^{\text{p}}$ can be obtained by the low-complexity one-dimensional search..
@@ III-B3 Outer layer iteration
In the outer layer, we initialise a large $\varrho$ and update $\varrho$ at each outer iteration by $\varrho=\varrho\bar{c}$, where $10<\bar{c}<$ 1 is the iteration coefficient of the penalty terms. The penalty-based two-layer algorithm is summarized in Algorithm 1.
@@ III-C Overall Algorithm
Algorithm 1 Penalty-based two-layer algorithm.
1:  Initialize $\mathbf{x}^{\text{p}}$, $\mathbf{u}_{i}$, and $\mathbf{v}_{i}$. Set the convergence accuracy $\epsilon_{1}$, $\epsilon_{2}$, and $\epsilon_{3}$.
@@
3:      update $\{\mathbf{W},\mathbf{R_{\text{s}}}\}$ by solving the problem (P4).
@@
The proposed penalty-based AO algorithm is summarized in Algorithm 2, which is assured to converge at least to a stationary point solution. The computational complexity of Algorithm 2 mainly depends on solving the SDP problems (P4), (P7), and the one-dimensional exhaustive search. It is given by $\mathcal{O}\Big{(}\log(\frac{1}{\epsilon_{4}})((N_{1})^{3.5}+\log(\frac{1}{\epsilon_{3}})\log(\frac{1}{\epsilon_{2}})\big{[}\log(\frac{1}{\epsilon_{1}})(N_{1})^{3.5}+N\bar{Q}\big{]})\Big{)}$ [8], where $\bar{Q}$ represents the number of the quantization bits during the one-dimensional exhaustive search.
@@ IV Numerical Results
Figure 3: The equivalent illumination power versus the transmit power at the BS.
Figure 4: The equivalent illumination power versus the rotation angle of the dielectric waveguide..
Figure 5: The equivalent illumination power versus the rotation angle of the dielectric waveguide.
This section evaluates the performance of the proposed PASS-ISAC framework. A 3D topological network setup is considered, where the dielectric waveguide is located in the x-o-z plane with a height of $d$ and a length of $50$ m. The communicating user and the sensing target are located in a square region centered at the origin in the x-o-y plane. The default simulation parameters are set as: $\sigma^{2}=-105$ dBm, $f=28$ GHz, $d=10$ m, $\Delta x=\frac{\lambda}{2}$, $r_{\text{s}}=30$ m, $\phi_{\text{s}}=\frac{\pi}{3}$, $[x_{1},y_{1},z_{1}]=[-15,-15,0]$ m, $N=12$, $\eta_{\text{eff}}=1.4$, $R_{\text{QoS}}=2$ bps/Hz, $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=10^{-3}$. The other network parameters are shown in the captions of the figures.
@@
  • Fixed antenna: In this scheme, $N$ pinching antennas are uniformly spread along the dielectric waveguide.
  • MRC scheme: In the maximum-ratio combining (MRC) scheme, the communication and sensing beamformer are designed by $\mathbf{w}=\frac{\sqrt{P_{\text{max}}}}{2}\frac{\mathbf{h}_{1,\text{c}}}{\|\mathbf{h}_{1,\text{c}}\|}$ and $\mathbf{s}=\frac{\sqrt{P_{\text{max}}}}{2}\frac{\mathbf{h}_{2,\text{s}}}{\|\mathbf{h}_{s,\text{s}}\|}$ [4].
  • ZF scheme: In the zero-forcing (ZF) scheme, the communication and sensing beamformer are designed by $[\mathbf{w},\mathbf{s}]=\mathbf{H}(\mathbf{H}^{H}\mathbf{H})^{-1}\mathbf{P}$, where $\mathbf{H}=[\mathbf{h}_{1,\text{c}},\mathbf{h}_{2,\text{s}}]$ and $\mathbf{P}=\text{diag}([\frac{\sqrt{P_{\text{max}}}}{2},\frac{\sqrt{P_{\text{max}}}}{2}])$ [4].
Here, the equivalent illumination power is defined as the $\frac{P(\theta_{\text{s}},r_{\text{s}})}{\eta}$, which neglects the impact of the free-space path-loss coefficient $\eta$ on the illumination power. In Fig. 5, we can observe that the pinching antenna achieves the highest illumination power compared to the other baseline schemes. This result can be expected because compared with the conventional antenna and fixed antenna, pinching antennas can be flexibly repositioned to attenuate the large-scale path loss between the pinching antennas and the receiving ends. It indicates that the pinching antenna is capable of providing more spatial degrees-of-freedom (DoFs) to favor the communication and sensing performance. On the other hand, as the MRC and ZF schemes cannot guarantee the optimal beamforming design under the given antenna activation positions, they exhibit inferior performance to the pinching antenna scheme.
Fig. 5 depicts the relationship between the equivalent illumination power and the number of activated pinching antennas, with a comparison of the optimal power allocation model (i.e., each pinching antenna can radiate an arbitrary proportion of the entire transmit power) and proportional power allocation $(P_{n}=\delta(\sqrt{1-\delta^{2}})^{n-1}P_{\text{max}})$ models [10] is considered. Without loss of generality, $\delta$ is set as $0.5$. As can be observed, the equal power allocation model achieves a comparable performance to that of the considered ideal power allocation model. This is because, in the equal power model, each pinching antenna will be allocated a constant level of power, which ensures the spatial DoF gain of the pinching antennas. However, in the proportional power model, pinching antennas located far away from the BS cannot obtain enough power to serve the nearby users, which limits the number of effective pinching antennas. We also observe that increasing $r_{\text{s}}$ decreases the illumination power at the target due to a significant increase in large-scale path loss.
@@ V Conclusion
A novel PASS-ISAC framework has been proposed, where the transmitting array was evenly split into two parts for supporting independent communication and sensing beamforming designs. A penalty-based AO algorithm was proposed to maximize the illumination power at the target while guaranteeing the QoS requirement of the communication user. Simulation results were provided to verify the superiority of the proposed PASS-ISAC framework over the other benchmarks.

### v2

@@ III-B Antenna Activation Position Optimization
To deal with the intractable expression of variable channels, we introduce auxiliary variables $\mathbf{U}$ and $\mathbf{V}$ to replace $\mathbf{H}_{\text{s}}$ and $\mathbf{H}_{\text{c}}$, respectively. Thus, we have the equality constraints $\mathbf{U}=\mathbf{H}_{\text{s}}$ and $\mathbf{V}=\mathbf{H}_{\text{c}}$. By relocating the equality constraint to the objective function and serving as a penalty term, the problem (P3) can be equivalently rewritten as
$(\text{P}5)\quad$ $\max\limits_{\mathbf{x}^{\text{p}},\mathbf{U},\mathbf{V}}\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi$ (13a)
s.t. $\text{Tr}(\mathbf{V}\mathbf{W})\geq\gamma_{\text{QoS}}\sigma^{2},$ (13b)
@@
$\frac{1}{r_{\text{min,s}}^{2}}\leq\frac{\mathbf{U}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,s}}^{2}},$ (13e)
$\frac{1}{r_{\text{min,c}}^{2}}\leq\frac{\mathbf{V}_{[n,n]}}{\eta}\leq\frac{1}{r_{\text{max,c}}^{2}},$ (13f)
@@
where $\chi=\|\mathbf{U}-\mathbf{H}_{\text{s}}\|_{F}+\|\mathbf{V}-\mathbf{H}_{\text{c}}\|_{F}$, and $\varrho$ denotes the scaling factor of the penalty terms. Here, $r_{\text{min,s}}$ and $r_{\text{max,s}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the sensing target. Also, $r_{\text{min,c}}$ and $r_{\text{max,c}}$ denote the minimum and maximum distances from an arbitrary pinching antenna to the communication user. To solve the problem (P5), we propose a penalty-based two-layer algorithm, which alternately optimizes $\{\mathbf{U},\mathbf{V}\}$ and $\{\mathbf{x}^{\text{p}}\}$ in the inner layer and update $\varrho$ in the outer layer.
@@ III-B1 Inner layer iteration—subproblem with respect to $\{\mathbf{U},\mathbf{V}\}$
$(\text{P}6)\ \max\limits_{\mathbf{U},\mathbf{V}}$ $\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi$ (14a)
@@
$\begin{cases}\Re(\text{Tr}(\mathbf{U}^{H}(\mathbf{I}-\mathbf{u}\mathbf{u}^{H})))\leq\varrho_{1},\\ \Re(\text{Tr}(\mathbf{V}^{H}(\mathbf{I}-\mathbf{v}\mathbf{v}^{H})))\leq\varrho_{2},\end{cases}\ i\in\{1,2\},$ (15)
where $\mathbf{u}$ and $\mathbf{v}$ represent the leading eigenvectors of $\mathbf{U}$ and $\mathbf{V}$. As a result, the problem (P6) can be transformed into
$(\text{P}7)\ \max\limits_{\mathbf{U},\mathbf{V},\varrho_{i}}\quad$ $\text{Tr}(\mathbf{U}\mathbf{W})-\frac{1}{2\varrho}\chi-\sum_{i=1}^{2}\varrho_{i}$ (16a)
@@
which is a convex problem and can be directly solved. Thus, the rank-one solution $\{\mathbf{U},\mathbf{V}\}$ can be obtained by iteratively solving the problem (P7).
@@ III-B2 Inner layer iteration—subproblem with respect to $\{\mathbf{x}^{\text{p}}\}$
Note that the equality constraint $\mathbf{U}=\mathbf{H}_{\text{s}}$ and $\mathbf{V}=\mathbf{H}_{\text{c}}$ are equivalent to $\mathbf{u}=\mathbf{h}_{\text{s}}$ and $\mathbf{v}=\mathbf{h}_{\text{c}}$, where $\{\mathbf{u},\mathbf{v}\}$ denote the leading eigenvectors of $\{\mathbf{U},\mathbf{V}\}$. As a result, the problem (P6) can be transformed into
$(\text{P}8)\ \min\limits_{\mathbf{x}^{\text{p}}}\quad$ $\|\mathbf{u}-\mathbf{h}_{\text{s}}\|^{2}+\|\mathbf{v}-\mathbf{h}_{\text{c}}\|^{2}$ (17a)
@@
It is easy to notice that $x_{n}^{\text{p}}$ and $x_{m}^{\text{p}}$ ($n\neq m$) are separated in the objective function but coupled in the constraint (8f), which motivates us to adopt the element-wise optimization framework. Therefore, with the fixed $\{x_{1}^{\text{p}},\cdots,x_{n-1}^{\text{p}},x_{n+1}^{\text{p}},\cdots,x_{N}^{\text{p}}\}$, the subproblem with respect to $x_{n}^{\text{p}}$ is given by
@@
$\varpi(x_{n}^{\text{p}})=\|\mathbf{u}_{[n]}-\mathbf{h}_{[n],\text{s}}\|^{2}+\|\mathbf{v}_{[n]}-\mathbf{h}_{[n],\text{c}}\|^{2}.$ (19)
Then, the optimal $x_{n}^{\text{p}}$ can be obtained by the low-complexity one-dimensional search.
@@ III-B3 Outer layer iteration
In the outer layer, we initialise a large $\varrho$ and update $\varrho$ at each outer iteration by
$\varrho=\varrho\bar{c},$ (20)
where $10<\bar{c}<$ 1 is the iteration coefficient of the penalty terms. The two-layer algorithm is summarized in Algorithm 1.
@@ III-C Overall Algorithm
Algorithm 1 Two-layer algorithm.
1:  Initialize $\mathbf{x}^{\text{p}}$, $\mathbf{u}$, and $\mathbf{v}$. Set the convergence accuracy $\epsilon_{1}$, $\epsilon_{2}$, and $\epsilon_{3}$.
@@
3:      update $\{\mathbf{W}\}$ by solving the problem (P4).
@@
The proposed penalty-based AO algorithm is summarized in Algorithm 2, which is assured to converge at least to a stationary point solution. The computational complexity of Algorithm 2 mainly depends on solving the SDP problems (P4), (P7), and the one-dimensional exhaustive search. It is given by $\mathcal{O}\Big{(}\log(\frac{1}{\epsilon_{4}})\left(N^{3.5}+\log(\frac{1}{\epsilon_{3}})\log(\frac{1}{\epsilon_{2}})\big{[}\log(\frac{1}{\epsilon_{1}})N^{3.5}+N\bar{Q}\big{]}\right)\Big{)}$ [8], where $\bar{Q}$ represents the number of the quantization bits during the one-dimensional exhaustive search.
@@ IV Numerical Results
Figure 3: The illumination power versus the transmit power at the BS.
Figure 4: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{max}}=70$ dBm.
Figure 5: The illumination power versus the rotation angle of the dielectric waveguide, where $P_{\text{max}}=70$ dBm.
This section evaluates the performance of the proposed PASS-ISAC framework. A 3D topological network setup is considered, where the dielectric waveguide is located in the x-o-z plane with a height of $d$ and a length of $50$ m. The communicating user and the sensing target are located in a square region centered at the origin in the x-o-y plane. The default simulation parameters are set as: $\sigma^{2}=-105$ dBm, $f=28$ GHz, $d=10$ m, $\Delta x=\frac{\lambda}{2}$, $r_{\text{s}}=30$ m, $\phi_{\text{s}}=\frac{\pi}{3}$, $[x_{1},y_{1},z_{1}]=[-15,-15,0]$ m, $N=16$, $\eta_{\text{eff}}=1.4$, $R_{\text{QoS}}=10$ bps/Hz, $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=\epsilon_{4}=10^{-3}$. The other network parameters are shown in the captions of the figures.
@@
  • Fixed pinching antenna: In this scheme, $N$ pinching antennas are uniformly spread along the dielectric waveguide.
  • Low-complexity MRC-based scheme: In the low-complexity maximum-ratio combining (MRC)-based scheme, $\mathbf{w}$ is designed by aligning with both $\mathbf{h}_{\text{c}}$ and $\mathbf{h}_{\text{s}}$ simultaneously, i.e., $\mathbf{w}=\sqrt{\frac{P_{\text{max}}}{N}}(\mathbf{h}_{\text{c}}+\mathbf{h}_{\text{s}})\odot\frac{1}{\mathbf{\check{h}}}$, where $\mathbf{\check{h}}=[|\mathbf{h}_{[1],\text{c}}+\mathbf{h}_{[1],\text{s}}|,\cdots,|\mathbf{h}_{[N],\text{c}}+\mathbf{h}_{[N],\text{s}}|]$.
In Fig. 5, we can observe that the pinching antenna achieves the highest illumination power compared to the other baseline schemes. This result can be expected because compared with the conventional antenna and fixed pinching antenna, pinching antennas can be flexibly repositioned to attenuate the large-scale path loss between the pinching antennas and the receiving ends. It indicates that the pinching antenna is capable of providing more spatial degrees-of-freedom (DoFs) to favor the communication and sensing performance. On the other hand, as the low-complexity MRC-based scheme cannot guarantee the optimal radiation pattern design under the given antenna activation positions, it exhibits inferior performance to the proposed scheme.
Fig. 5 depicts the relationship between the illumination power and the number of activated pinching antennas, with a comparison of an upper bound (i.e., each pinching antenna can radiate an arbitrary proportion of the entire transmit power) and proportional power allocation $(P_{n}=\delta(\sqrt{1-\delta^{2}})^{n-1}P_{\text{max}})$ model [10]. Without loss of generality, $\delta$ is set as $0.5$. As can be observed, the equal power allocation model achieves a comparable performance to that of the upper bound. This is because, in the equal power model, each pinching antenna will be allocated a constant level of power, which ensures the spatial DoF gain of the pinching antennas. However, in the proportional power model, pinching antennas located far away from the BS cannot obtain enough power to serve the nearby users, which limits the number of effective pinching antennas. We also observe that increasing $r_{\text{s}}$ decreases the illumination power at the target due to a significant increase in large-scale path loss.
@@ V Conclusion
A novel PASS-ISAC framework has been proposed, where the pinching beamforming was exploited to realize the simultaneous C&S transmission. A separated ISAC design was proposed for the two-waveguide PASS. A penalty-based AO algorithm was proposed to maximize the illumination power at the target while guaranteeing the QoS requirement of the communication user. Simulation results were provided to verify the superiority of the proposed PASS-ISAC framework over the other benchmarks.